Publications of James Edward McDonald

Institute of Atmospheric Physics, University of Arizona

(compiled by K.T. McDonald, updated March 31, 2025)

Click on the blue links to download .pdf copies.

J.E. McDonald’s typewritten version of his publication list:

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_publications.pdf

Publications in Refereed Journals

1. (with K.E. Penrod and G.L. Maison) The Rate of Descent of Parachutes from Various Altitudes, J. Aeronaut. Sci. 14, 303-310 (1947),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_14_303_47.pdf

2. Remarks on “Sea Salt in a Tropical Storm”, J. Meteor. 8, 362 (1951),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_8_362_51.pdf

3. Erroneous Cloud-Physics Applications of Raoult’s Law, J. Meteor. 10, 68 (1953),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_10_68_53.pdf

4. Remarks on “Atmospheric salt particles and raindrops”, J. Meteor. 10, 165 (1953),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_10_165_53.pdf

5. Homogeneous Nuclear of Supercooled Water Drops, J. Meteor. 10, 416 (1953),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_10_416_53.pdf

6. Remarks on the Development of the Coriolis Principle, Bull. Am. Meteor. Soc. 34,

  192 (1953), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_34_192_53.pdf

7. On the Electrical Conductivity of the Lower Stratosphere, Trans. Am. Geophys. Un.

  34, 529 (1953), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tagu_34_529_53.pdf

8. (with J.P. Lodge and F. Baer) An Investigation of the Melander Effect, J. Meteor.

  11, 318 (1954), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_11_318_54.pdf

9. The Shape and Aerodynamics of Large Raindrops, J. Meteor. 11, 478 (1954),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_11_478_54.pdf

10. Reply to Mason, J. Meteor. 11, 515 (1954),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_11_515_54.pdf

11. Scientific Basis of Weather Modification, Science 124, 86 (1956),

   http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_124_86_56.pdf

12. A note on the precision of estimation of missing precipitation data, Trans. Am. Geophys. Union. 38, 657 (1957),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tagu_38_657_57.pdf

13. “Deposition” — a proposed antonym for “sublimation”, J. Meteor. 15, 245 (1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_15_245_58.pdf

also in Am. J. Phys. 26, 131 (1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_26_131_58.pdf

also in J. Chem. Phys. 28, 170 (1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jcp_28_170_58.pdf

14. Nomenclature of Phase Transition, J. Chem. Ed. 35, 205 (1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jce_35_205_58.pdf

also in Quart. J. Roy. Meteor. Soc. 84, 74 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_qjrms_84_74_58.pdf

15. Mean Upper-Air Data for Selected World Stations, Mon. Weath. Rev. 87, 299 (1959),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_mwr_87_299_59.pdf

16. An Analysis of Effects of Nuclear Attacks on Tucson and Phoenix, J. Ariz. Acad. Sci.

   1, 41 (1959), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jaas_1_41_59.pdf

17. Cloud Photography and Cloud Physics Research, Weather 14, 247 (1959),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weather_14_247_59.pdf

18. Variability Factors in Mountain-Watershed Hydrometeorology in an Arid Region, J. Ariz. Acad. Sci. 1, 89 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jaas_1_89_60.pdf

19. Absorption of Atmospheric Radiation by Water Films and Water Clouds, J. Meteor.

   17, 232 (1960), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_17_232_60.pdf

20. Direct Absorption of Solar Radiation by Atmospheric Water Vapor, J. Meteor. 17,

   319 (1960), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_17_319_60.pdf

21. Rates of descent of fallout particles from thermonuclear explosions, J. Meteor. 17,

   380 (1960), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_17_380_60.pdf

22. (with G.G. Goyer, F. Baer and R.R. Braham, Jr) Effects of Electric Fields on Water-Droplet Coalescence, J. Meteor. 17, 442 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_17_442_60.pdf

23. An Aid to Computation of Terminal Fall Velocities of Spheres, J. Meteor. 17, 463

   (1960), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_17_463_60.pdf

24. A Note on the Aerodynamic Effects of Internal Circulations in Small Raindrops, J. Meteor. 17, 474 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_17_474_60.pdf

25. (with C.R. Green) A Comparison of Rank-Difference and Product-Moment Correlation of Precipitation Data, J. Geophys. Res. 65, 333 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jgr_65_333_60.pdf

26. (with C.R. Green) Effects of Inhomogeneity and Record Length on Estimates of Correlation and Variability of Precipitation Data, J. Geophys. Res. 65, 2375 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jgr_65_2375_60.pdf

27. Remarks on Correlation Methods in Geophysics, Tellus 12, 176-183 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tellus_12_176_60.pdf

28. Comments on “An Airborne Cloud-Drop-Size Distribution Meter”, J. Meteor. 18,

   423 (1961), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_18_423_61.pdf

29. On Radar Detection of Non-spherical Ice Particles, J. Meteor. 18, 695 (1961),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jm_18_695_61.pdf

30. An Analysis of Civil Defense Hazards Being Created by Emplacement of Intercontinental Ballistic Missiles near Tucson, J. Ariz. Acad. Sci. 2, 3 (1961),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jaas_2_3_61.pdf

31. What is a “Meteorological Engineer”?, Bull. Am. Meteor. Soc. 42, 122 (1961),

     http://http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_122_61.pdf

32. On the Ratio of Evaporation to Precipitation, Bull. Am. Meteor. Soc. 42, 185 (1961),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_185_61.pdf

33. An Historical Note on an Early Cloud-Modification Experiment, Bull. Am. Meteor. Soc. 42 195 (1961),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_195_61.pdf

34. Mean Atmospheric Residence Times for Particulate Air Pollutants, Bull. Am. Meteor. Soc. 42 664 (1961),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_664_61.pdf

35. The Surface Chemistry of Ice Nucleation, Bull. Am. Meteor. Soc. 42, 723-725 (1961),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_723_61.pdf

36. Airlight Simulation Experiments, J. Atmos. Sci. 19, 114 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_19_114_62.pdf

37. (with H.D. Orville) Effect of Momentum-Exchange on the Fall Velocity of an Accreting Hydrometer, J. Atmos. Sci. 19, 278 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_19_278_62.pdf

38. Large-sphere limit of the radar back-scattering coefficient, Quart. J. Roy. Meteor. Soc. 88, 183-186 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_qjrms_88_183_62.pdf

39. The evaporation-precipitation fallacy, Weather 17, 168-177 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weather_17_168_62.pdf

40. A gigantic horizontal cloudbow, Weather 17, 243-245 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weather_17_243_62.pdf

41. Collection and Washout of Airborne Pollens and Spores by Raindrops, Science 135,

   435 (1962), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_135_435_62.pdf

42. A Note on Anomalous Adiabatic Cooling Rates in Clouds, J. Atmos. Sci. 19, 309

   (1962), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_19_309_62.pdf

43. Visibility Reduction due to Jet-Exhaust Carbon Particle, J. Appl. Meteor. 1, 391

   (1962), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jam_1_391_62.pdf

44. Homogenous Nucleation of Vapor Condensation. I. Thermodynamic Aspects, Am. J. Phys. 30, 870 (1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_30_870_62.pdf

45. Some Models of Interest in Meteorological Thermodynamics, Bull. Am. Meteor. Soc.

   43, 292 (1962), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_43_292_62.pdf

46. Intermolecular Attractions and Saturation Vapor Pressure, J. Atmos. Sci. 20, 178

   (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_20_178_63.pdf

47. Accuracy of Numerical Evaluations of the Pseudoadiabatic Cooling Rate, J. Atmos.

   Sci. 20, 240 (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_20_240_63.pdf

48. The Saturation Adjustment in Numerical Modeling of Fog, J. Atmos. Sci. 20, 476

   (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_20_476_63.pdf

49. Stratospheric Cloud over Northern Arizona, Science 140, 292 (1963),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_140_292_63.pdf

50. Status of Investigation of the Northern Arizona Stratospheric Cloud of February 28,

   1963, http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_630228.pdf

51. Homogenous Nucleation of Vapor Condensation. II. Kinematic Aspects, Am. J. Phys.

   31, 31 (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_31_31_63.pdf

52. Caustic of the Primary Rainbow, Am. J. Phys. 31, 282 (1963),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_31_282_63.pdf

53. Rain washout of partially wettable insoluble particles, J. Geophys. Res. 68, 4993

   (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jgr_68_4993_63.pdf

54. Use of the Electrostatic Analogy in Studies of Ice Crystal Growth, Z. Angew. Math.

   Phys. 14, 610 (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_zamp_14_610_63.pdf

55. Early Developments in the Theory of the Saturated Adiabatic Process, Bull. Am. Meteor. Soc. 44, 203-211 (1963),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_44_203_63.pdf

56. James Espy and the Beginnings of Cloud Thermodynamics, Bull. Am. Meteor. Soc.

   44, 634-641 (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_44_634_63.pdf

57. Meridional Alignment of Tropical Cirrus Clouds, Nature 201, 1316 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nature_201_1316_64.pdf

58. On a Criterion Governing the Mode of Cloud Formation in Planetary Atmospheres, J. Atmos. Sci. 21, 76 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_21_76_64.pdf

59. Cloud Nucleation on Insoluble Particles, J. Atmos. Sci. 21, 109 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_21_109_64.pdf

60. A Thermodynamic Relation in the Theory of Homogenous Nucleation of Supercooled Droplets, J. Atmos. Sci. 21, 225 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jas_21_225_64.pdf

61. Pollen Wettability as a Factor in Washout by Raindrops, Science 143, 1180 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_143_1180_64.pdf

62. (with B.M. Herman) On a point of terminology in the optics of the rainbow, Bull. Am. Meteor. Soc. 45, 279-280 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_45_279_64.pdf

63. Atmospheric exclusion limits for clouds of water and other substances, J. Geophys. Res. 69, 3669 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jgr_69_3669_64.pdf

64. Raindrop Spatter-Spots, Weather 19, 177-179 (1964),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weather_19_177_64.pdf

65. Reply to Moyer, J. Geophys. Res. 70, 248 (1965),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jgr_70_248_65.pdf

66. Saturation Vapor Pressures over Supercooled Water, J. Geophys. Res. 70, 1553

   (1965), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_jgr_70_1553_65.pdf

67. Maxwellian Interpretation of the Laplacian, Am. J. Phys. 33, 706 (1965),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_33_706_65.pdf

Chapters of Books

68. Paleoclimatology, Chap. 12 of Geochronology, U. Ariz. Phys. Sci. Bull. Series 26, No. 2, p. 196 (April 1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_uapsb_26_2_196_55.pdf

69. The Physics of Cloud Modification, Adv. in Geophys. 5, 233-303 (1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ag_5_233_58.pdf

70. Evaluation of Weather Modification Field Tests, in Weather Modification, Science and Public Policy, ed. by R.G. Fleagle (U. Wash. Press, 1969),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_wmspp_69.pdf

Publications in Popular Journals

71. The Coriolis Effect, Sci. Am. 186, 72-76 (May 1952),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_sa_186_5_72_52.pdf

reprinted in New Worlds of Modern Science, ed. by L. Engel (Dell, 1956),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nwms_88_56.pdf

72. Letter, Sky and Telescope 11, 158 (May 1952),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_st_11_158_52.pdf

73. The Earth’s Electricity, Sci. Am. 188, 32-37 (Apr. 1953),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_sa_188_4_32_53.pdf

74. The Shape of Raindrops, Sci. Am. 190, 64-68 (Feb. 1954),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_sa_190_2_64_54.pdf

75. Collected Fragments of a Whilom Aeropoet, Weather 12, 107 (1957),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weather_12_107_57.pdf

76. The Great Aurora of September 22-23, Sky and Telescope 17, 14 (Nov. 1957),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_st_17_11_14_57.pdf

77. Letter, Sky and Telescope 18, 179 (Feb. 1958),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_st_18_2_179_58.pdf

78. “It rained everywhere but here! The thunderstorm encirclement illusion, Weatherwise 174, 158 (Aug. 1959),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weatherwise_12_158_59.pdf

79. (with C.E. Brooks and M.G Wurtele) Contribution to The Glossary of Meteorology, ed. by R.E. Husckle (American Meteorological Society, Boston, 1959).

80. Cities into Targets, The Nation 190, 436 (1960),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nation_190_436_60.pdf

81. The Ice-Fall Problem, Weatherwise 13, 110 (June 1960),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weatherwise_13_110_60.pdf

82. Should We Know All the Facts about Civil Defense?, Point West 4, 24 (Feb. 1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_pw_4_2_24_62.pdf

83. An Aid to Study of the Heavens, Sky and Telescope 22, 134 (Mar. 1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_st_22_134_62.pdf

84. Tucson and the Titans, Arizona Frontiers 1, 9 (Nov. 1962),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_af_1_5_9_61.pdf

85. Cloud-Ring in the Upper Stratosphere, Weatherwise 16, 99 (June 1963),

    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weatherwise_16_99_63.pdf

86. The Big City Atom, Frontier 14, 5 (Sept. 1963),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_frontier_14_9_5_63.pdf

Letters to Editors

87. Annual and Seasonal Persistence of Precipitation in Arizona, Bull. Am. Meteor. Soc. 41, 106 (1960),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_41_106_60.pdf

88. “Backlash”, Science 133, 1271 (1961),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_133_1271_61.pdf

89. The Evaporation-Precipitation Fallacy, Weather 17, 216 (1962),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_weather_17_216_62.pdf

90. Jets, Meteors, Missiles, or Rockets?, Science 141, 1136 (1963),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_141_1136_63.pdf

91. Atmospheric Jet Streams, Science 143, 489 (1964),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_143_489_64.pdf

92. Moondoggle, Science 144, 1086 (1964),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_144_1086_64.pdf

93. Evolution in Arizona, Science 151, 632 (1966),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_151_632_66.pdf

Book Reviews

94. Physical Meteorology, by J.C. Johnson, Bull. Am. Meteor. Soc. 36, 40 (1955),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_36_40_55.pdf

95. A Concise Dictionary of Science, by F. Gaynor, Bull. Am. Meteor. Soc. 41, 93

(1960), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_41_93_60.pdf

96. The Mechanics of Aerosols, by N.A. Fuchs, Bull. Am. Meteor. Soc. 42, 221 (1961),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_221_61.pdf

97. Fluids Dynamic Drag, by S.F. Hoerner, Bull. Am. Meteor. Soc. 42, 302 (1961),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_42_302_61.pdf

98. Aerodynamic Capture of Particles, by E.G. Richardson, Bull. Am. Meteor. Soc. 43,

113 (1962), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_43_113_62.pdf

99. Thinking about the Unthinkable, by H. Kahn, The Progressive, 26, 44 (Oct. 1962),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_progressive_26_44_oct_62.pdf

100. Kill and Overkill, by R.E. Lapp, The Nation 195, 379 (1962),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nation_195_379_62.pdf

101. Evaporation and Droplet Growth in Gaseous Media, by N.A. Fuchs, Bull. Am. Meteor. Soc. 44, 103 (1963),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_44_103_63.pdf

102. The Microbiology of the Atmosphere, by P.H. Gregory, Bull. Am. Meteor. Soc. 44,

267 (1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_44_267_63.pdf

103. Clouds, Rain and Rainmaking, by B.J. Mason, Trans. Am. Geophys. Un. 44, 153

(1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tagu_44_154_53.pdf

104. Air Chemistry and Radioactivity, by C.E. Junge, Trans. Am. Geophys. Un. 44, 1019

(1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tagu_44_1019_63.pdf

105. Strategy for Survival, by T.L. Martin and D.C. Latham, The Progressive, 27, 46

(Oct. 1963), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_progressive_27_46_oct_63.pdf

106. The Atomic Age, ed. by M. Grodzins and E. Rabinowitch, The Progressive, 28, 40

(Apr. 1964), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_progressive_28_10_40_64.pdf

107. An Introduction to Fluid Mechanics and Heat Transfer, by J.M. Kay, Am. J. Phys.

32, 389 (1964), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_32_389_64.pdf

108. Mathematics and the Physical World, by M. Kline, Am. J. Phys. 32, 716 (1964),

http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ajp_32_716_64.pdf

109. An Introduction to Atmospheric Physics, by R.G. Fleagle and J.A. Businger, Trans. Am. Geophys. Un. 45, 323 (1964),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tagu_45_323_64.pdf

110. Environmental Radioactivity, by M. Eisenbud, Trans. Am. Geophys. Un. 45, 456

    (1964), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_tagu_45_456_64.pdf

111. The Physics of Rainclouds, by N.H. Fletcher, Science 143, 236 (1964),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_143_236_64.pdf

112. Aerodynamics, by J.E. Allen, Science 144, 707 (1964),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_144_707_64.pdf

113. The Mechanics of Aerosols, by N.A. Fuchs, Science 146, 1033 (1964),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_146_1033_64.pdf

114. Weather, by P.D. Thompson, R. O’Brien, and the Editors of Life, and Weather Eyes in the Sky: America’s Meteorological Satellites by J.G. Vaeth, Science 149, 739

    (1965), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_149_739_65.pdf

115. Problems of Atmospheric and Space Electricity, ed. by S.C. Coroniti, Science 149,

    1085 (1965), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_science_149_1085_65.pdf

116. Thermodynamics of Clouds, by L. Dufour and R. Defay, Bull. Am. Meteor. Soc. 46,

    674 (1965), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_46_674_65.pdf

117. Physics of the Earth’s Upper Atmosphere, ed. by C.O. Hines et al., Bull. Am. Meteor. Soc. 46, 803 (1965),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bams_46_803_65.pdf

118. Cosmic Rays, by B. Rossi, Am. Sci. 53, 380A (1965),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_as_53_380a_65.pdf

Conference Contributions

119. Measurement, Analysis, and Interpretation of Turbulent Temperature Fluctuations, Int. Symp. Atmos. Turb. Bound. Layer (MIT, June 1951),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_grp_19_263_51.pdf

120. (with A.R. Kassander and L.L. Sims) Observations of Freezing Nuclei over the Southwestern U.S., Proc. Conf. Cloud Phys. Precip. Part. (Woods Hole, 1955),

    p. 392, http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ccppp_392_55.pdf

121. Increasing Rain Artificially, Proc. Conf. Water for Texas 2, 23 (1956),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_pcwt_2_23_56.pdf

Technical Reports

122. Summer Air Mass Instability and Its Synoptic Representation, Master’s Thesis (MIT,

    Feb. 12, 1945), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_masters_thesis_45.pdf

123. Radiation-Errors and Lag-Errors in the Measurement of Turbulent Temperature Fluctuations, Ph.D. Thesis (Iowa State University, 1952),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_phd_thesis_51.pdf

124. (with A.H. Woodcock) Sea Salt Nuclei Studies, Woods Hole Oceanographic Inst. Rep. 52-75 (Sep. 1952),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_whoi52-75.pdf

125. Theoretical Cloud Physics Studies, Office of Naval Research project 082083 (Jan. 13,

    1953), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_onr082093_53.pdf

126. (with A.R. Kassander) Lag Effects in the Measurement of Turbulent Temperature Fluctuations, Iowa State College Dept. Phys. Rep. No. 1 (Feb. 1, 1952).

127. (with A.R. Kassander and L.L. Sims) Observations of Freezing Nuclei over the Southwestern U.S., Inst. Atmos. Phys. Sci. Rep. No. 3 (Nov. 1,1956),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_iap3.pdf

128. Variability of Precipitation in an Arid region: A Survey of Characteristics for Arizona, Inst. Atm. Phys. Tech. Rep. No. 1 (Dec. 31, 1956),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_iap1.pdf

129. A Critical Evaluation of Correlation Methods in Climatology and Hydrology, Inst. Atmos. Phys. Sci. Rep. No. 4 (Jan. 25, 1957),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_iap4.pdf

130. Cloudiness over the Southwestern United States and Its Relation to Astronomical Observing, Inst. Atmos. Phys. Sci. Rep. No. 7 (Feb. 24, 1958),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_iap7.pdf

131. Meteorological Aspects of Nuclear Reactor Hazards at Bodega Bay, (June, 1964),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_bodega_bay.pdf

132. Relationship of skin cancer incidence to thickness of ozone layer, statement submitted to hearings before the House Subcommittee on Transportation Appropriations (Mar. 2, 1971), Cong. Rec. 117, 7252 (1971),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_cr_117_7251_71.pdf

133. A Variational Derivation of Young’s Equation for the Contact Angle (draft, Mar. 22,

    1971), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_032271.pdf

Publications and Statements on Unidentified Flying Objects

134. Talks on Anomalous Aerial Phenomena (Oct. 1966 – Mar. 1968),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_talks_66-68.pdf

135. The Problem of Unidentified Flying Objects, summary of a talk presented to the Am. Meteor. Soc. (Washington, DC, Oct. 19, 1966),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_ams_101966.pdf

136. Comments on Australian radio regarding his tour of Australia (Feb. 1967),

     http://kirkmcd.princeton.edu/JEMcDonald/australia_feb67.mp4

137. UFOs: Greatest Scientific Problem of Our Times?, presented at the annual meeting of the Am. Soc. Newspaper Eds. (Washington, DC, Apr. 22, 1967),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_asne_67.pdf

138. International Scientific Aspects of the Problem of Unidentified Flying Objects, presented to the Group of Outer Space Affairs of the United Nations (New York,

    June 7, 1967), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_un_060767.pdf

139. Unidentified Flying Objects: Greatest Scientific Problem of Our Times, coempendium of several papers by NICAP (1967),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nicap_67.pdf

140. UFOs: Extraterrestrial Probes? Aeronautics and Astronautics 5(8), 19 (1967),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_aa_5_8_19_67.pdf

141. Science, Technology, and UFOs, presented at a General Seminar of the United Aircraft Research Laboratories (East Hartford, CT, Jan. 26, 1968),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_uarl_012668.pdf

142. UFOs – An International Scientific Problem, Astronautics Symposium, Canadian Aeronautics and Space Institute (Montreal, Mar. 12, 1968),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_031268.pdf

143. Are UFOs Extraterrestrial Surveillance Craft? American Institute of Aeronautics and Astronautics (Los Angeles, Mar. 26, 1968),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_aiaa_032668.pdf

144. A Need for an International Study of UFOs, Flying Saucer Review (London) 14(2),

    11 (1968), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_fsr_14_2_11_68.pdf

145. Statement on Unidentified Flying Objects, submitted to House Committee on Science and Astronautics (Washington, DC, July 29, 1968),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_hcsa_68.pdf
     http://files.ncas.org/ufosymposium/mcdonald.html
     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_680729.pdf

146. UFOs and the Condon Report, prepared for the Medical Students Colloquium of the Arizona College of Medicine (Tucson, AZ, Jan. 30, 1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_acm_013069.pdf

147. UFOs and the Condon Report – A Scientist’s Critique, presented to the Dupont Chapter of The Scientific Research Society of America (RESA, Wilmington, DE, Feb.

    12, 1969), http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_resa_021269.pdf

148. UFOs – A Challenge to Observation, presented to the Symposium on Meteorological Observations and Instrumentation (Washington, DC, Feb. 13, 1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_smoi_021369.pdf

149. UFOs and the Condon Report: A Dissenting View, presented at the Pacific Missile Range Section, Am. Inst. Aero. Astro. (Pt. Mugu, CA, Feb. 18, 1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_021869.pdf

150. The Dissenting View, Engineering Opportunites (April, 33, 1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_eo_april_33_69.pdf

151. UFOs – Unsolved: A Scientific Challenge, NICAP (Washington, DC, June, 10, 1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nicap_061069.pdf

152. Review of The Condon Report, Scientific Study of Unidentified Flying Objects, Icarus 11, 443 (1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_icarus_11_443_69.pdf

153. Science in Default: Twenty-Two Years of Inadequate UFO Investigations, presented at the 136^{th} meeting of the Am. Assoc. Adv. Sci. (Boston, Dec. 27, 1969),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_aaas_69.pdf
     http://kirkmcd.princeton.edu/JEMcDonald/science_in_default_aass69.mp4

154. UFOs over Lakenheath in 1956, Flying Saucer Review (London) 16(2), 9 (1970),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_fsr_16_2_9_70.pdf

155. The 1957 Gulf Coast B-47 Incident, Flying Saucer Review (London) 16(3), 2 (1970),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_fsr_16_3_2_70.pdf

156. The Kirtland Airfield UFO, Flying Saucer Review (London) 16(5), 6 (1970),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_fsr_16_5_6_70.pdf

157. Meteorological Factors in Unidentified Radar Returns, presented at the 14^{th} Radar Meteor. Conf. (Tucson, Nov. 1970),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_rmc_1170.pdf
    http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_amsrms_70.pdf (in French!)

158. Letter to the Hon. Silvio Conte (Mar. 8, 1971),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_030871.pdf

159. UFO Encounter 1. Sample Case Selected by the UFO Subcommittee of the AIAA, Aeronautics and Astronautics 9(7), 66 (1971),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_aa_9_7_66_71.pdf

160. Papers on Unidentified Flying Objects and Related Phenomena (1958-1971), University of Arizona Special Collection MS412 (previously MS328)

      http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_arizona_archives_Ms328.pdf
      http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_arizona_archives_ms412.pdf
        http://sabio.library.arizona.edu.ezproxy1.library.arizona.edu/search~S9/X?SEARCH=%28James%20McDonald%29&searchscope=9&SORT=D&m=p&b=sc
      http://www.azarchivesonline.org/xtf/view?docId=ead/uoa/UAMS412.xml;query=;brand=default

161. ETH: The ExtraTerrestrial Hypothesis, Audio recording (date uncertain),

    WorldCat 129644203, http://www.worldcat.org/

162. Quotations assembled over many years,

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_quotes.pdf
163. Archive of audio recordings, http://wolfdigitalmedia.com/ufo/jamesmcdonaldarchive.htm

Publications about James McDonald

Some newspaper articles about James McDonald are at

     http://kirkmcd.princeton.edu/Family/James

164. James McDonald, A Cloud Physicist (NY Times, July 7, 1971),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_nytimes_710716.pdf

165. L.J. Battan, A.R. Kassander, Jr, P.S. Martin and C. Steelink, Memorial Resolution: James Edward McDonald (1971),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_memorial.pdf

166. L.J. Battan, Obituary of James E. McDonald, Bull. Am. Meteor. Soc. 52, 735 (1971),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_obituary_bams_71.pdf

167. P.E. McCarthy, Politicking and Paradigm Shifting: James E. McDonald and the UFO Case Study, Ph.D. disseration (U. Hawaii, Dec. 1975),

     http://files.ncas.org/mccarthy/index.html

168. UFOs and Science: The Collected Writings of Dr. James E. McDonald, ed. by

    V. Vaughn (One Reed Publications, 1995), http://www.onereed.com/books/ufobook.php

169. A. Druffel, Firestorm: Dr. James E. McDonald’s Fight for UFO Science, 2nd ed. (Granite Publishing, 2003),

     http://kirkmcd.princeton.edu/JEMcDonald/druffel_firestorm.pdf
     http://kirkmcd.princeton.edu/JEMcDonald/xcon_2004.mp4
170. Wikipedia biography: http://en.wikipedia.org/wiki/James_E._McDonald

171. Some aspects of James E. McDonald’s UFO research can be traced at the Project

    Bluebook website: http://www.bluebookarchive.org/default.aspx

172. D. Ouse, Forgotten Duluthian: James E. McDonald, UFO Researcher, Zenith City online (Dec. 6, 2012),

     http://kirkmcd.princeton.edu/JEMcDonald/mcdonald_zenith_city_120612.pdf
     http://zenithcity.com/forgotten-duluthian-james-e-mcdonald-ufo-researcher/

173. J.K. Gillin, Looking Up: Professor James E. McDonald 1955-1970, Senior Thesis (U. Arizona, June 2013),

     http://kirkmcd.princeton.edu/JEMcDonald/gillin_senior_thesis_jem_0613.pdf

174. K. Reagan, UA Professor Created Own ‘X-Files’, Arizona Public Media online (July

    3, 2013), http://kirkmcd.princeton.edu/JEMcDonald/apm_070313.pdf
      https://originals.azpm.org/p/uaplus-stories/2013/7/3/25213-ua-professor-created-own-x-files/
     http://kirkmcd.princeton.edu/JEMcDonald/ufo_files.mp4

175. A. Foster, Audio reveals creepy details of UFO mystery, news.com.au (Aug. 8, 2018),

     http://kirkmcd.princeton.edu/JEMcDonald/foster_news.com.au_080818.pdf
     http://kirkmcd.princeton.edu/JEMcDonald/westall_ufo.mp4